Because 10% of the reservation holders are 'no shows' a Canadian airline sells 400 tickets for a flight that can accommodate 375 passengers.
a) Find the probability that one or more reservation holders will not be accomodated on the flight.
b) Find the probability of fewer than 350 passengers on the flight.

Answer by stanbon(75887)

Because 10% of the reservation holders are 'no shows' a Canadian airline sells 400 tickets for a flight that can accommodate 375 passengers.
a) Find the probability that one or more reservation holders will not be accomodated on the flight.
---
Binomial Problem with n = 400, p = 0.1
P(x>375) = 1 - P(0<= x <=374) = 1 - 400C374(0.9)^374*(0.1)^26 = 0.0054...
-----------------------------------------
b) Find the probability of fewer than 350 passengers on the flight.
P(0<= x <=349) = 0.0436..
-----------------------------------------
Cheers,
Stan H.